OpenStudy (anonymous):

if you add 5 to me and then subtract 2, you get 4

OpenStudy (anonymous):

\[x=me\] haha \[(x+5)-2=4\] \[x+3=4\] \[x+3-3=4-3\] \[x+0=1\] \[x=1\] So I am 1

OpenStudy (anonymous):

you started with 3

OpenStudy (anonymous):

its 1 since 1 + 5 = 6 but you subtract 2, 6 - 2 = 4

OpenStudy (anonymous):

what is the actual question